# The Multi-Wavelength Context of Delayed Radio Emission in TDEs: Evidence for Accretion-Driven Outflows

### v1

@@ Abstract
Recent observations presented in Cendes et al. (2024a) show that optically selected tidal disruption events (TDEs) commonly produce delayed radio emission that can peak years post-disruption. Here, we explore the multi-wavelength properties of a sample of radio-observed optically selected TDEs to shed light on the physical process(es) responsible for the late-rising radio emission. We combine new late-time X-ray observations with archival optical, UV, X-ray, and radio data to conclude that a diversity of accretion-driven outflows may power delayed radio emission in TDEs. Our analysis suggests that some late radio outflows may be launched by a delayed phase of super-Eddington accretion onto the central supermassive black hole (SMBH), while others may result from a state transition to a “low-hard” radiatively inefficient accretion flow or the deceleration of an off-axis relativistic jet. We additionally find weak statistical evidence that TDEs with delayed radio emission have larger optical/UV photospheric radii than other TDEs and are less likely to exhibit helium emission lines at early times, possibly also supporting the hypothesis that the onset of SMBH accretion is delayed in these systems. Our results have implications for our understanding of state changes in SMBH accretion flows, the circularization timescale for TDE debris, and the prevalence of off-axis jets in TDEs, and motivates systematic, long-term monitoring of these unique transients. The brightest objects in our sample are also detected in the VLA Sky Survey (VLASS), demonstrating that all-sky radio surveys can play an important role in discovering unexpected radio properties of the TDE population.
@@
facilities: VLA, MeerKAT, ATCA, Swift(XRT and UVOT), Chandra, XMM-Newtonsoftware: astropy (Astropy Collaboration et al., 2013, 2018), mosfit (Guillochon et al., 2018), Source Extractor (Bertin & Arnouts, 1996)
1 Introduction
@@
The strong gravitational field near a supermassive black hole (SMBH) is capable of tidally disrupting an approaching star, temporarily enhancing the mass fallback rate onto the SMBH by orders of magnitude as gas from the star is accreted (Hills, 1975; Rees, 1988). Non-thermal synchrotron emission from these tidal disruption events (TDEs) is best observed in the radio or millimeter and has been variously ascribed to jets, non-relativistic outflows, interactions between unbound stellar debris and the circumnuclear medium, or some combination thereof (see Alexander et al. 2020 for a review). A small fraction ($\lesssim 1$%) of known TDEs produce powerful relativistic jets with peak radio luminosities $\nu L_{\nu}\gtrsim 10^{41}$ erg s-1 (Zauderer et al., 2011; Cenko et al., 2012; Brown et al., 2015; Andreoni et al., 2022), but most TDEs lack such jets and are instead discovered as long-lasting transients in X-ray or optical surveys (e.g., Bade et al. 1996; Saxton et al. 2021; Sazonov et al. 2021; van Velzen et al. 2021; Hammerstein et al. 2023a; Yao et al. 2023). Targeted radio follow up of individual X-ray and optically discovered TDEs has revealed that some such TDEs exhibit lower-luminosity radio emission ($\nu L_{\nu}\lesssim 10^{39}$ erg s-1) within a few months of discovery (e.g., van Velzen et al. 2016; Alexander et al. 2016, 2017; Stein et al. 2021; Cendes et al. 2021a), but early follow up efforts typically ceased if no radio emission was detected within $6-12$ months, leaving the late-time radio properties of most TDEs unknown.
@@
The first searches for TDE radio emission on longer timescales of months to years post-discovery were optimized to target initially-relativistic jets viewed off-axis, as the radio emission of such jets would be suppressed at early times due to relativistic beaming and only become visible after jet deceleration. These searches resulted in only non-detections, confirming that the rarity of powerful relativistic jets was a physical and not a selection effect; however, these studies were not sensitive enough to probe the existence of weaker outflows on similar timescales (Komossa, 2002; van Velzen et al., 2013; Bower et al., 2013). Recently, deeper wide-field radio transient searches have resulted in the first blindly selected radio TDE candidates (i.e., nuclear radio transients discovered independently, without prior detection of a counterpart at other wavelengths), suggesting that fainter radio emission in TDEs may be $\sim 10$ times more common than powerful jets (Anderson et al. 2020; see also Somalwar et al. 2023a; Dykaar et al. 2024). However, it is still not clear if these radio-selected nuclear transients probe the same underlying population as optically selected or X-ray selected TDEs.
@@
This is the second of two papers comprising the first large-scale attempt to quantify the prevalence of fainter, late-rising radio emission in TDEs and understand the physical origin of this emission. The first of the two papers, Cendes et al. (2024a) (hereafter C24), presented the surprising result that late-rising radio emission is common in optically selected TDEs, with $\sim 40$% of the observed objects showing variable radio emission years post-discovery. This is comparable to — or possibly slightly higher than — the rate of TDEs that exhibit “prompt” radio emission (i.e., within 6 months of discovery), although the sample size of TDEs with early radio observations remains small and the occurrence rate of prompt radio emission in TDEs thus remains poorly constrained (Alexander et al., 2020). The delayed radio flares characterized by C24 and others (e.g., Horesh et al. 2021a, b; Cendes et al. 2022b; Hajela et al. 2024) are generally faint ($\nu L_{\nu}\lesssim 10^{39}$ erg s-1, with a few exceptions) and have been variously ascribed to non-relativistic outflows launched several hundred days post-disruption (Horesh et al., 2021a; Cendes et al., 2022b), to off-axis relativistic jets launched promptly (Matsumoto & Piran, 2023; Sfaradi et al., 2024), to initially-choked precessing relativistic jets that only escape after an initially misaligned accretion disk aligns with the black hole spin axis (Teboul & Metzger, 2023; Lu et al., 2024), or to outflowing material encountering changes in the density of the ambient circumnuclear medium (Matsumoto & Piran, 2024; Zhuang et al., 2024). While the physical origin of TDE radio emission remains unclear, the wide range of emission onset times — from a few days post-discovery to $>1000$ days post-discovery — may suggest that multiple mechanisms are responsible for producing low-luminosity radio emission in TDEs.
@@
The multi-wavelength properties of TDEs may shed light on the origin of late-rising radio emission, as they give us independent methods to trace the evolution of the SMBH accretion rate with time. Mapping the accretion rate has been attempted in a variety of ways, most commonly by modeling TDEs’ optical/UV or X-ray light curves (e.g. Mockler et al. 2019; Saxton et al. 2021). If delayed radio brightenings are caused by newly-launched accretion-driven outflows, then we expect some correlation between the time when radio emission first appears and the instantaneous accretion rate onto the SMBH at that time. Conversely, if the delayed radio brightenings are powered by a promptly-launched outflow encountering changes in the ambient environment, we would not expect such correlations.
@@
While the optical through X-ray emission from TDEs is generally broadly consistent with thermal emission from hot gas near the SMBH, the exact connection between the observed optical, UV, and X-ray light curves of TDEs and the accretion rate has been the subject of intense theoretical debate (Roth et al. 2020 and references therein). One open question is whether the mass fallback rate (i.e. the rate at which stellar debris approaches the black hole from the point of disruption) maps directly to the accretion rate (i.e. the rate at which material crosses the event horizon), or if the latter is significantly delayed relative to the former. While the mass fallback rate may be easier to constrain observationally in some cases, it is the accretion rate that is relevant for predicting the launch times of radio-producing outflows. For this work, we therefore consider two extremal possibilities: 1) that the stellar debris rapidly circularizes and forms an accretion flow, such that the bolometric optical-through-UV light curve directly traces the mass fallback rate (typically expected to exceed the Eddington limit before declining as $t^{-5/3}$; Rees 1988) and the instantaneous accretion rate onto the SMBH (accounting for a minimal viscous delay, e.g., Guillochon & Ramirez-Ruiz 2013; Mockler et al. 2019); or 2) that the peak accretion rate onto the SMBH is significantly delayed relative to peak optical light, either because the debris circularization is inefficient (in which case the optical peak may be powered instead by shocks from stream-stream collisions, e.g., Piran et al. 2015; Shiokawa et al. 2015; Lu & Bonnerot 2020) or because radiation pressure initially prevents even rapidly circularized debris from collapsing into a disk while the mass fallback rate is high, instead forming a hot, extended quasi-spherical pressure-supported envelope that powers the early optical emission while keeping the accretion rate across the event horizon low (Loeb & Ulmer, 1997; Guillochon et al., 2014; Metzger, 2022). In Case 1, X-rays produced near the SMBH are reprocessed into the optical band by more extended material, so the accretion rate is best traced in the optical/UV (Coughlin & Begelman, 2014; Guillochon et al., 2014; Metzger & Stone, 2016; Roth et al., 2016; Dai et al., 2018; Mockler et al., 2019). In Case 2, tracking the X-ray luminosity may provide a better method of estimating the instantaneous accretion rate, as thermal X-ray emission in these models originates in the innermost part of the accretion flow and thus indicates that matter is successfully approaching the SMBH regardless of the origin of the optical emission (Auchettl et al., 2017; Gezari et al., 2017; Metzger, 2022). In both cases, state transitions in the accretion flow are expected as the accretion rate changes, causing the production and quenching of outflows, but the timing of these transitions relative to peak optical light will be different. We compare the onset of any observed radio emission to the expected timing of state transitions in the accretion flow in both cases, focusing on the highly super-Eddington and highly sub-Eddington accretion phases that are most likely to power outflows (e.g., Giannios & Metzger 2011; De Colle et al. 2012).
@@
In C24, we modeled the radio emission detected in optically selected TDEs at late times using an equipartition analysis, discussed its properties in the context of other radio observations of TDEs, and calculated the occurrence rate of late-time radio emission. Here, we discuss the multi-wavelength properties of radio-observed optical TDEs in more detail and attempt to discriminate among possible physical models for delayed radio emission. In Section 2 we introduce our sample and describe our new multi-wavelength observations and the archival data that we use for our subsequent analysis. We describe our light curve modeling and compare the radio properties of our sample to their multi-wavelength emission and derived properties in Section 3. We discuss the implications of our results in Section 4 and we conclude in Section 5.
2 Sample definition and data collection
@@
This work was originally motivated by a search for radio emission associated with 91 known TDE candidates in the VLA Sky Survey (VLASS). This search resulted in three detections of transient radio emission associated with high-confidence TDEs and three detections of likely unrelated radio sources (Appendix A). Surprisingly, the detected TDEs included a dramatic rebrightening of the optical TDE ASASSN-15oi discovered in VLASS epoch 1 at $\sim 3$ years post-disruption (Horesh et al., 2021a; Hajela et al., 2024), motivating the deeper radio searches presented in C24 and this paper. A second optical TDE in our sample, AT2018hyz, appears as an even brighter radio transient in VLASS epoch 3 (Cendes et al. 2022b; C24).
@@
To probe the physical mechanism or mechanisms behind the newly discovered late-rising radio emission in optically detected TDEs, we utilize multi-wavelength observations of a sample of TDEs discovered in optical surveys between 2014 and 2020 (Table 1). Our sample contains 31 TDEs at redshift $z\lesssim 0.1$ for which targeted radio follow up was obtained. As we wish to understand if late-time radio flares have a different physical origin than prompt radio emission in TDEs, our sample consists of the 24 TDEs observed by C24, most of which had no published radio detections prior to that work, plus an additional 7 TDEs from the literature previously reported to exhibit radio emission at early times. In addition to compiling optical, UV, and X-ray photometry of our TDEs from the literature, we also obtained new late-time X-ray observations of all TDEs in our sample. We briefly summarize all data used in this paper below and provide details of our new data reduction and analysis. A complete list of all references for previously-published data used in this work is given in Appendix B (Table 4), along with all new radio and X-ray photometry used for our analysis (Tables 5, 6). The complete multiwavelength dataset will also be made available as part of the Open mulTiwavelength Transient Event Repository (OTTER; Franz et al. in prep).
@@
Table 1: Our Sample: Optically-selected TDEs at $z\lesssim 0.1$ with radio observations discovered between January 2014 and October 2020. TDEs in bold are those with newly discovered radio emission, presented in full in our companion paper, C24.
@@
These TDEs showed only H lines in their initial classification spectra, leading to their labeling as TDE-H in some references. However, here we classify them as TDE-H+He because they also exhibit He emission at $\gtrsim 50-200$ days.
iPTF16fnl has been previously referred to in the literature as a “delayed” radio TDE due to several radio upper limits at early times (Horesh et al., 2021b). However, its first radio detection is $153$ $(151)$ days post-discovery in the observer (rest) frame, which places it in the “prompt” category as defined here (Section 2.1).
2.1 Radio: VLA, MeerKAT, ATCA, and literature data
In C24, we presented new radio observations at 2 – 6 years post-discovery of 24 optically-selected TDEs at $z\lesssim 0.1$ discovered between January 2014 and October 2020. These 24 TDEs are drawn from the samples defined by van Velzen et al. (2021) and Hammerstein et al. (2023a), plus a few additional events from the literature (Foley et al., 2015; Blanchard et al., 2017; Gromadzki et al., 2019; Frederick et al., 2020; Angus et al., 2022). Following C24, we also include previously-published radio observations and public archival data from the VLASS and VAST radio surveys (Lacy et al., 2020; Murphy et al., 2013) when available. For this work, we reanalyzed one observation of PS16dtm previously reported in C24, taken with NSF’s Karl G. Jansky Very Large Array (the VLA) on 2017 Aug 22 (372 days post-discovery) under program 16B-318 (PI: Alexander), and report a tentative new detection of $39\pm 10$ $\mu$Jy at 6 GHz. For AT2018fyk, we also include the recent radio detections reported by Cendes et al. (2024b).
In this paper, we add 7 TDEs from the literature with promptly-detected radio emission, to compare and contrast with the late-rising radio TDEs. Six of these (ASASSN-14li, ASASSN-15oi, AT2019azh, ASASSN-19bt, AT2019qiz, and AT2020opy) are included in the population analyses by van Velzen et al. (2021) and/or Hammerstein et al. (2023a); the remaining TDE is the youngest object in our sample and the first TDE classified as part of the second phase of the Zwicky Transient Facility survey (ZTF-II; Bellm et al. 2019a), AT2020vwl (Yao et al., 2023; Goodwin et al., 2023a). For AT2019qiz, we also include a subset of the VLA observations collected under programs 19A-013, 20A-372 (PI: Alexander), and 21A-303 (PI: Hajela), a public observation collected as part of the VLASS epoch 3, and a MeerKAT observation collected under program SCI-20220822-YC-01 (PI: Cendes). The AT2019qiz radio data used in our analysis are listed in Table 5 in Appendix B, and the complete radio dataset for AT2019qiz will be presented in future work (Franz et al. in prep).

### v2

@@ Abstract
Recent observations presented in Cendes et al. (2024a) show that optically selected tidal disruption events (TDEs) commonly produce delayed radio emission that can peak years post-disruption. Here, we explore the multi-wavelength properties of a sample of radio-observed optically selected TDEs to shed light on the physical process(es) responsible for the late-rising radio emission. We combine new late-time X-ray observations with archival optical, UV, X-ray, and radio data to conclude that a diversity of accretion-driven outflows may power delayed radio emission in TDEs. Our analysis suggests that some late radio outflows may be launched by a delayed phase of super-Eddington accretion onto the central supermassive black hole (SMBH), while others may result from a state transition to a “low-hard” radiatively inefficient accretion flow or the deceleration of an off-axis relativistic jet. We find that TDEs with delayed radio emission are less likely to exhibit helium emission lines at early times ($p=0.002$) and may have larger optical/UV photospheric radii ($p=0.026$) than other TDEs, possibly also indicating that the onset of SMBH accretion is delayed in these systems. Our results have implications for our understanding of state changes in SMBH accretion flows, the circularization timescale for TDE debris, and the prevalence of off-axis jets in TDEs, and motivate systematic, long-term monitoring of these unique transients. The objects in our sample with the brightest radio emission are also detected in the VLA Sky Survey (VLASS), demonstrating that all-sky radio surveys can play an important role in discovering unexpected properties of the TDE population.
@@
facilities: VLA, MeerKAT, ATCA, Swift(XRT and UVOT), Chandra, XMM-Newtonsoftware: astropy (Astropy Collaboration et al., 2013, 2018), mosfit (Guillochon et al., 2018), Source Extractor (Bertin and Arnouts, 1996)
I Introduction
@@
The strong gravitational field near a supermassive black hole (SMBH) is capable of tidally disrupting an approaching star, temporarily enhancing the mass fallback rate onto the SMBH by orders of magnitude as gas from the star is accreted (Hills, 1975; Rees, 1988). Non-thermal synchrotron emission from these tidal disruption events (TDEs) is best observed in the radio or millimeter and has been variously ascribed to jets, non-relativistic outflows, interactions between unbound stellar debris and the circumnuclear medium, or some combination thereof (see Alexander et al. 2020 for a review). A small fraction ($\lesssim 1$%) of known TDEs produce powerful relativistic jets with peak radio luminosities $\nu L_{\nu}\gtrsim 10^{41}$ erg s-1 (Zauderer et al., 2011; Cenko et al., 2012; Brown et al., 2015; Andreoni et al., 2022), but most TDEs lack such jets and are instead discovered as long-lasting transients in X-ray or optical surveys (e.g., Bade et al. 1996; Saxton et al. 2021; Sazonov et al. 2021; van Velzen et al. 2021; Hammerstein et al. 2023b; Yao et al. 2023). Targeted radio follow up of individual X-ray and optically discovered TDEs has revealed that some such TDEs exhibit lower-luminosity radio emission ($\nu L_{\nu}\lesssim 10^{39}$ erg s-1) within a few months of discovery (e.g., van Velzen et al. 2016; Alexander et al. 2016, 2017; Stein et al. 2021; Cendes et al. 2021a), but early follow up efforts typically ceased if no radio emission was detected within 6126-12 months, leaving the late-time radio properties of most TDEs unknown.
@@
The first searches for TDE radio emission on longer timescales of months to years post-discovery were optimized to target initially-relativistic jets viewed off-axis, as the radio emission of such jets would be suppressed at early times due to relativistic beaming and only become visible after jet deceleration. These searches resulted in only non-detections, confirming that the rarity of powerful relativistic jets was a physical and not a selection effect; however, these studies were not sensitive enough to probe the existence of weaker outflows on similar timescales (Komossa, 2002; van Velzen et al., 2013; Bower et al., 2013). Recently, deeper wide-field radio transient searches have resulted in the first blindly selected radio TDE candidates (i.e., nuclear radio transients discovered independently, without prior detection of a counterpart at other wavelengths), suggesting that fainter radio emission in TDEs may be $\sim 10$ times more common than powerful jets (Anderson et al. 2020; see also Somalwar et al. 2025; Dykaar et al. 2024). However, it is still not clear if these radio-selected nuclear transients probe the same underlying population as optically selected or X-ray selected TDEs.
@@
This is the second of two papers comprising the first large-scale attempt to quantify the prevalence of less luminous, late-rising radio emission in TDEs and understand the physical origin of this emission. The first of the two papers, Cendes et al. (2024a) (hereafter C24), presented the surprising result that late-rising radio emission is common in optically selected TDEs, with $\sim 40$% of the observed objects showing variable radio emission years post-discovery. This is comparable to — or possibly slightly higher than — the rate of TDEs that exhibit “prompt” radio emission (i.e., within 6 months of discovery), although the sample size of TDEs with early radio observations remains small and the occurrence rate of prompt radio emission in TDEs thus remains poorly constrained (Alexander et al., 2020). The delayed radio flares characterized by C24 and others (e.g., Horesh et al. 2021a, b; Cendes et al. 2022a; Hajela et al. 2025) are generally faint ($\nu L_{\nu}\lesssim 10^{39}$ erg s-1, with a few exceptions) and have been variously ascribed to non-relativistic outflows launched several hundred days post-disruption (Horesh et al., 2021a; Cendes et al., 2022a), to off-axis relativistic jets launched promptly (Matsumoto and Piran, 2023; Sfaradi et al., 2024), to initially-choked precessing relativistic jets that only escape after an initially misaligned accretion disk aligns with the black hole spin axis (Teboul and Metzger, 2023; Lu et al., 2024), or to outflowing material encountering changes in the density of the ambient circumnuclear medium (Matsumoto and Piran, 2024; Zhuang et al., 2025). While the physical origin of TDE radio emission remains unclear, the wide range of emission onset times — from a few days post-discovery to >1000>1000 days post-discovery — may suggest that multiple mechanisms are responsible for producing low-luminosity radio emission in TDEs.
@@
The multi-wavelength properties of TDEs may shed light on the origin of late-rising radio emission, as they give us independent methods to trace the evolution of the SMBH accretion rate with time. Mapping the accretion rate has been attempted in a variety of ways, most commonly by modeling TDEs’ optical/UV or X-ray light curves (e.g., Mockler et al. 2019; Saxton et al. 2021). If delayed radio brightenings are caused by newly-launched accretion-driven outflows, then we expect some correlation between the time when radio emission first appears and the instantaneous accretion rate onto the SMBH at that time. Conversely, if the delayed radio brightenings are powered by a promptly-launched outflow encountering changes in the ambient environment, we would not expect such correlations.
@@
While the optical through X-ray emission from TDEs is generally broadly consistent with thermal emission from hot gas near the SMBH, the exact connection between the observed optical, UV, and X-ray light curves of TDEs and the accretion rate has been the subject of intense theoretical debate (e.g., Piran et al. 2015; Shiokawa et al. 2015; Roth et al. 2016; Dai et al. 2018; Mockler et al. 2019; Lu and Bonnerot 2020; Roth et al. 2020 and references therein; Ryu et al. 2020a; Metzger 2022; Sarin and Metzger 2024). One open question is whether the mass fallback rate $\dot{M}_{\rm fb}$ (i.e., the rate at which stellar debris approaches the black hole from the point of disruption) maps directly to the accretion rate $\dot{M}$ (i.e., the rate at which material crosses the event horizon), or if the latter is significantly delayed relative to the former. The mass fallback rate may be easier to constrain observationally in some cases, as it is generally assumed to be directly proportional to the bolometric luminosity, $L=\epsilon\dot{M}_{\rm fb}c^{2}$, where $\epsilon$ is a constant (typically $\sim 0.1$). However, it is the accretion rate that is relevant for predicting the launch times of radio-producing outflows. Measuring the accretion rate is less straightforward due to the inherent messiness of the disruption process and debris circularization. For this work, we therefore consider two extremal possibilities; namely Case 1: that the stellar debris rapidly circularizes and forms an accretion flow, such that the bolometric optical-through-UV light curve directly traces both the mass fallback rate (typically expected to exceed the Eddington limit before declining as $t^{-5/3}$; Rees 1988) and the instantaneous accretion rate onto the SMBH (accounting for a minimal viscous delay, e.g., Guillochon and Ramirez-Ruiz 2013; Mockler et al. 2019); or Case 2: that the peak accretion rate onto the SMBH is significantly delayed relative to peak optical light, either because the debris circularization is inefficient (in which case the optical peak may be powered instead by shocks from stream-stream collisions, e.g., Piran et al. 2015; Shiokawa et al. 2015; Lu and Bonnerot 2020) or because radiation pressure initially prevents even rapidly circularized debris from collapsing into a disk while the mass fallback rate is high, instead forming a hot, extended quasi-spherical pressure-supported envelope that powers the early optical emission while keeping the accretion rate across the event horizon low (Loeb and Ulmer, 1997; Guillochon et al., 2014; Metzger, 2022). In Case 1, X-rays produced near the SMBH are reprocessed into the optical band by more extended material, so the accretion rate is best traced in the optical/UV and can be computed by fitting the optical/UV light curves with existing modeling codes (Coughlin and Begelman, 2014; Guillochon et al., 2014; Metzger and Stone, 2016; Roth et al., 2016; Dai et al., 2018; Mockler et al., 2019). In Case 2, tracking the X-ray luminosity and comparing it to model predictions may provide a better method of estimating the instantaneous accretion rate, as thermal X-ray emission in these models originates in the innermost part of the accretion flow and thus indicates that matter is successfully approaching the SMBH regardless of the origin of the optical emission (Auchettl et al. 2017; Gezari et al. 2017; Metzger 2022; Sarin and Metzger 2024). In both cases, state transitions in the accretion flow are expected as the accretion rate changes, causing the production and quenching of outflows, but the timing of these transitions relative to peak optical light will be different. We compare the onset of any observed radio emission to the expected timing of state transitions in the accretion flow in both cases, focusing on the highly super-Eddington and highly sub-Eddington accretion phases that are most likely to power outflows (e.g., Giannios and Metzger 2011; De Colle et al. 2012).
@@
C24 modeled the radio emission detected in optically selected TDEs at late times using an equipartition analysis, discussed its properties in the context of other radio observations of TDEs, and calculated the occurrence rate of late-time radio emission. Here, we discuss the multi-wavelength properties of radio-observed optical TDEs in more detail and attempt to discriminate among possible physical models for delayed radio emission. In Section II we introduce our sample and describe our new multi-wavelength observations and the archival data that we use for our subsequent analysis. We describe our light curve modeling and compare the radio properties of our sample to their multi-wavelength emission and derived properties in Section III. We discuss the implications of our results in Section IV and we conclude in Section V.
II Sample definition and data collection
@@
This work was originally motivated by a search for radio emission associated with 91 known TDE candidates in the VLA Sky Survey (VLASS). This search resulted in three detections of transient radio emission associated with high-confidence TDEs and three detections of likely unrelated radio sources (Appendix A). Surprisingly, the detected TDEs included a dramatic rebrightening of the optical TDE ASASSN-15oi discovered in VLASS epoch 1 at $\sim 3$ years post-disruption (Horesh et al., 2021a; Hajela et al., 2025), motivating the deeper radio searches presented in C24 and this paper. A second optical TDE in our sample, AT2018hyz, appears as an even brighter radio transient in VLASS epoch 3 (Cendes et al. 2022a; C24).
@@
To probe the physical mechanism or mechanisms behind the newly discovered late-rising radio emission in optically detected TDEs, we utilize multi-wavelength observations of a sample of TDEs discovered in optical surveys between 2014 and 2020 (Table 1). Our sample contains 31 TDEs for which targeted radio follow up was obtained. Most are at redshift $z\lesssim 0.1$; we also include two TDEs at slightly higher $z$ with constraining early radio data (DES14C1kia and AT2020opy). As we wish to understand if late-time radio flares have a different physical origin than prompt radio emission in TDEs, our sample consists of the 24 TDEs observed by C24, most of which had no published radio detections prior to that work, plus an additional 7 TDEs from the literature previously reported to exhibit radio emission at early times. In addition to compiling optical, UV, and X-ray photometry of our TDEs from the literature, we also obtained new late-time X-ray observations of all TDEs in our sample. We briefly summarize all data used in this paper below and provide details of our new data reduction and analysis. A complete list of all references for previously-published data used in this work is given in Appendix B (Table 6), along with all new radio and X-ray photometry used for our analysis (Tables 7, 8). The complete multiwavelength dataset will also be made available as part of the Open mulTiwavelength Transient Event Repository (OTTER; Franz et al. 2025).
@@
Table 1: Our Sample: Optically-selected TDEs at $z\lesssim 0.1$ with radio observations discovered between January 2014 and October 2020. TDEs in bold are those with newly discovered radio emission, presented in full in C24.
@@
C24 classify these TDEs as TDE-H+He because they exhibit He emission at $\gtrsim 50-200$ days. However, we follow other authors (e.g., van Velzen et al. 2021; Hammerstein et al. 2023a) in classifying them as TDE-H because they showed only H lines in their initial classification spectra around peak optical light.
iPTF16fnl has been previously referred to in the literature as a “delayed” radio TDE due to several radio upper limits at early times (Horesh et al., 2021b). However, its first radio detection is $153$ $(151)$ days post-discovery in the observer (rest) frame, which places it in the “prompt” category as defined here (Section II.1).
II.1 Radio: VLA, MeerKAT, ATCA, and literature data
[…]
In this paper, we add 7 TDEs from the literature with promptly-detected radio emission, to compare and contrast with the late-rising radio TDEs. Six of these (ASASSN-14li, ASASSN-15oi, AT2019azh, ASASSN-19bt, AT2019qiz, and AT2020opy) are included in the population analyses by van Velzen et al. (2021) and/or Hammerstein et al. (2023b); the remaining TDE is the youngest object in our sample and the first TDE classified as part of the second phase of the Zwicky Transient Facility survey (ZTF-II; Bellm et al. 2019b), AT2020vwl (Yao et al., 2023; Goodwin et al., 2023a). For AT2019qiz, we also include a subset of the VLA observations collected under programs 19A-013, 20A-372 (PI: Alexander), and 21A-303 (PI: Hajela), a public observation collected as part of the VLASS epoch 3, and a MeerKAT observation collected under program SCI-20220822-YC-01 (PI: Cendes). The AT2019qiz radio data used in our analysis are listed in Table 7 in Appendix B, and the complete radio dataset for AT2019qiz will be presented in future work (Franz et al. in prep).